40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 4 10:30:37 2016, ericq, Update, General, Preliminary vent plans 
    Reply  Tue Jul 5 17:14:02 2016, ericq, Update, General, Vent progress noOsemSide.JPGOSEMside.JPG
       Reply  Tue Jul 5 19:10:41 2016, rana, Update, General, Vent progress jumpEvent20151105.png
          Reply  Tue Jul 5 21:51:35 2016, ericq, Update, General, Vent progress 
             Reply  Tue Jul 5 23:38:42 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation image.jpeg
                Reply  Wed Jul 6 22:58:01 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                   Reply  Thu Jul 7 00:25:07 2016, ericq, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                      Reply  Thu Jul 7 10:49:03 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation UL_glitchMon_Striptool.png
                         Reply  Fri Jul 8 11:35:45 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation coilGlitchMon.png
                            Reply  Fri Jul 8 11:48:09 2016, Koji, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                               Reply  Fri Jul 8 14:03:26 2016, ericq, Update, General, ETMX removed from suspension zoom.jpg
                                  Reply  Fri Jul 8 16:33:26 2016, ericq, Update, General, ETMX removed from suspension 
                                     Reply  Fri Jul 8 20:44:09 2016, Koji, Update, General, ETMX removed from suspension 6x
                                        Reply  Sun Jul 10 15:38:01 2016, ericq, Update, General, ETMX in guiderod/standoff gluing fixture 6x
                                           Reply  Tue Jul 12 16:08:00 2016, ericq, Update, General, ETMX ruby guided gluing 
                                              Reply  Wed Jul 13 21:13:51 2016, Koji, Update, General, ETMX guide rod gluing / ETMY Magnet gluing 13x
                                                 Reply  Thu Jul 14 22:29:57 2016, ericq , Update, General, ETMX guide rod gluing / ETMY Magnet gluing 
                                                    Reply  Fri Jul 15 16:23:51 2016, ericq , Update, General, Gluing setbacks IMG_2768.JPGIMG_2771.JPG
                                                       Reply  Tue Jul 19 16:39:29 2016, Lydia, Update, General, ETMX magnet gluing/guiderod excess glue removed P7190201.JPGP7190203.JPG
                                                          Reply  Wed Jul 20 13:58:55 2016, Steve, Update, SUS, SUS damping out of vac chamber ETMXrack.jpgfromETMX-satBox.jpgsusDampingSatboxCab.jpgsusDampingSatboxCabl.jpg
                                                          Reply  Thu Jul 21 17:01:59 2016, ericq, Update, General, ETMX magnet gluing/guiderod excess glue removed 
                                                             Reply  Fri Jul 22 23:03:23 2016, gautam, Update, General, ETMX suspension - progress IMG_2828.JPGIMG_2826.JPGIMG_2821.JPGIMG_2822.JPGIMG_2823.JPG
                                                                Reply  Sun Jul 24 03:43:56 2016, ericq, Update, General, Magnet positioning 
                                                                Reply  Mon Jul 25 13:01:49 2016, ericq, Update, General, ETMX suspension - progress IMG_2829.JPGIMG_2830.JPGIMG_2831.JPGIMG_2832.JPGIMG_2833.JPG
                                                                   Reply  Mon Jul 25 13:47:19 2016, rana, Update, General, ETMX suspension - progress 
                                                                   Reply  Mon Jul 25 17:01:54 2016, Steve, Update, SUS, ETMX suspension - progress ITMYside2011.JPG
                                                                   Reply  Mon Jul 25 20:37:46 2016, gautam, Update, General, ETMX suspension - Side OSEM test osem_test_2016_07_25.png
                                                                      Reply  Tue Jul 26 18:07:57 2016, gautam, Update, General, ETMX suspension - Further OSEM Tests OSEM_spectra.pdf
                                                                         Reply  Thu Jul 28 03:22:28 2016, ericq, Update, General, ETMX magnets re-reglued 
                                                                            Reply  Thu Jul 28 13:49:44 2016, gautam, Update, General, ETMX magnets re-reglued ETMX_sideMagnet.JPGETMX_sideMagnet_zoomed.JPG
                                                                               Reply  Thu Jul 28 17:48:17 2016, gautam, Update, General, ETMY LR magnet broke off 
                                                                                  Reply  Thu Jul 28 22:43:03 2016, rana, Update, General, ETMY UL magnet broke off 
                                                                                  Reply  Wed Aug 3 10:52:03 2016, Steve, Update, General, ETMY soaking dish SD1.jpgSD2.jpgSD3.jpg
                                                                               Reply  Fri Jul 29 03:32:30 2016, ericq, Update, General, ETMX suspended once more 
                                                                                  Reply  Fri Jul 29 13:17:34 2016, Koji, Update, General, Oven 
                                              Reply  Thu Jul 14 01:00:33 2016, Koji, Update, General, ETMX ruby guided gluing ruby.jpg
                                  Reply  Thu Jul 28 16:43:01 2016, Lydia, Update, General, ETMX aluminium standoff groove condition wire_damage_4.jpgwire_damage_zoom.jpgno_wire.jpgcomparison_2.jpg
                            Reply  Fri Jul 8 21:22:38 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
             Reply  Wed Jul 6 09:34:26 2016, Steve, Update, General, PRM sensing voltage LR PRM_LR.png
             Reply  Wed Jul 6 19:36:45 2016, Koji, Update, General, SUS Vmon HPF.png160706_ITMX_VMON2.pdf160706_ITMX_VMON1.pdf
                Reply  Thu Jul 7 15:23:39 2016, ericq, Update, General, SUS Vmon coil_drive.png
                   Reply  Thu Jul 7 16:05:55 2016, Koji, Update, General, SUS Vmon 
       Reply  Thu Jul 7 17:12:50 2016, Johannes, Update, General, Vent progress IMG_2621.JPGIMG_2623.JPGIMG_2628.JPGIMG_2595.JPG
          Reply  Tue Jul 12 23:51:16 2016, Johannes, Update, General, Vent progress - ETMY inspection IMG_2679.JPGIMG_2673.JPG
             Reply  Wed Jul 13 00:38:25 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
                Reply  Tue Jul 19 13:21:42 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
       Reply  Sun Jul 10 17:33:00 2016, ericq, Update, General, Vent progress 
          Reply  Mon Jul 11 15:13:22 2016, gautam, Update, General, Vent progress: in-date First Contact procured image.jpeg
Message ID: 12335     Entry time: Mon Jul 25 20:37:46 2016     In reply to: 12331     Reply to this: 12340
Author: gautam 
Type: Update 
Category: General 
Subject: ETMX suspension - Side OSEM test 

[Lydia, gautam]

Summary: We did some preliminary tests to check if at least one of the side magnet positions is usable for the side OSEM. We mainly wanted to check how much dynamic range we lose because of the sub-optimal longitudinal positioning of the side magnet. We found that when the side magnet was mainly moving along the axis of the side OSEM (with minimal yaw motion as gauged by eye), the PD voltage bottomed out at ~80 counts (while the completely unoccluded readout was ~800 counts). 

Details:

  1. First, we placed the face OSEMs into their holders one by one, and adjusted their position till the readout was approximately half the saturation value (as judged by the average value of the readout, at this point, the mirror was still swinging around a fair bit).
  2. Next, we enabled the POS, PIT and YAW damping (with all existing settings unchanged), but with the SD coil input and output disabled. We had to increase the watchdog threshold to ~600mV.
  3. Once the optic was reasonably well damped (~70mV on the watchdog was the best we saw), I put in the side OSEM till the PD was completely occluded. At this point, I enganed the earthquake stops, and then released the mirror such that it was freely hanging. I then observed the optic by eye, and noted a time when the dominant motion was along the axis of the side OSEM coil (i.e. minimal YAW motion). 
  4. Attachment #1 shows time series plots of the 5 OSEM PD voltage monitors. Perhaps because the side OSEM input was disabled, the damping wasn't as efficient as it normally is (also there is a fan blowing air around the clean bench). But at the point indicated on the plot, the YAW motion was negligible to the eye, while the dominant motion was along the axis of the coil. During this time, the readout bottomed out at approximately 10% of the saturation value (towards the end of these plots, I disabled the damping loops and began pulling the OSEMs out one by one). Because the damping was imperfect, this is only an approximate guess of how much dynamic range we are losing. But does this warrant regluing the side magnet?

Other remarks:

  • The 4 face magents were reasonably well centered in the coil. While Eric and I were looking at this earlier today, the LL magnet looked a little close to the coil, but after putting all 4 OSEMs in, the situation looked reasonable to the eye. I couldn't take pictures because of space constraints, and furthermore, it's almost impossible to hold the camera in the correct vertical position. 
  • Steve, Eric and I couldn't find the OSEM gender changer anywhere in the lab and it wasn't in the box it was advertised to be in. So we made a custom cheater cable, and cleaned it by wiping with Isoprop., and wrapped it in foil for use in this test. The OSEM pins should probably be cleaned before we put these back in vacuum. 
Attachment 1: osem_test_2016_07_25.png  111 kB  | Show | Show all
ELOG V3.1.3-